# Scene-Adaptive Motion Planning with Explicit Mixture of Experts and Interaction-Oriented Optimization

### v1

@@ Appendix A Supplementary Material
In Figure 7, the scenario involves driving straight through an intersection. Here, Pluto unnecessarily changes lanes mid-intersection likely due to a perceived collision threat from the vehicle in the red box in Figure  (b), prompting a lane change that is reversed in Figure 7(c). EMOE-Planner, however, evaluates the collision potential more accurately, maintaining a straight trajectory and successfully navigating the intersection. However, although EMOE-Planner proceeds straight, the trajectory remains conservative due to the high traffic in the scenario. The current framework does not yet consider the prediction of other vehicles when planning the ego vehicle’s trajectory, which will be a focus of our future work.

### v2

@@ IV-D Simulation Results
In Figure 7, the scenario involves driving straight through an intersection. Here, Pluto unnecessarily changes lanes mid-intersection likely due to a perceived collision threat from the vehicle in the red box in Figure 7(b), prompting a lane change that is reversed in Figure 7(c). EMOE-Planner, however, evaluates the collision potential more accurately, maintaining a straight trajectory and successfully navigating the intersection. However, although EMOE-Planner proceeds straight, the trajectory remains conservative due to the high traffic in the scenario. The current framework does not yet consider the prediction of other vehicles when planning the ego vehicle’s trajectory, which will be a focus of our future work.
@@
V Conculsion
This paper introduces the EMOE-Planner, a novel framework tailored for enhancing trajectory planning in urban autonomous driving by utilizing scene information and interaction dynamics. The framework employs scene-specific queries generated through multiple scene anchors, focusing the model’s attention on relevant areas. Integration of a Mixture of Experts (MoE) with a shared scene router allows for precise matching of scene types with respective experts, improving the model’s responsiveness to varied scenarios and increasing the transparency of expert selection. Additionally, incorporating interactions between the ego vehicle and other agents into the loss function significantly improves trajectory scoring. Benchmark evaluations on several NuPlan datasets demonstrate that the EMOE-Planner surpasses state-of-the-art (SOTA) models by three points in closed-loop evaluations and closely aligns with the performance of post-processed SOTA methods, reducing reliance on rule-based post-processing. Furthermore, according to the simulation results, our model is the first pure learning model to achieve performance superior to rule-based algorithms (PDM-Closed) in NuPlan closed-loop simulations.
Acknowledgments
This work was supported by internal research funds from BYD Company Limited. No external funding was received. All authors are employees of BYD Company Limited. The research represents original work without conflicts of interest.